# Statistical description and dimension reduction of categorical trajectories with multivariate functional principal components

### v3

@@ Abstract
There are many examples in which the statistical units of interest are samples of a continuous time categorical random process, that is to say a continuous time stochastic process taking values in a finite state space. Getting simple representations that allow comparisons of a set of trajectories is of major interest for statisticians. Without loosing any information, we associate to each state a binary random function, taking values in $\{0,1\}$, and turn the problem of statistical description of the categorical trajectories into a multivariate functional principal components analysis. The (multivariate) covariance operator has nice interpretations in terms of departure from independence of the joint probabilities and the multivariate functional principal components are simple to interpret. Under the weak hypothesis assuming only continuity in probability of the 010-1 trajectories, it is simple to build consistent estimators of the covariance kernel and perform multivariate functional principal components analysis. The sample paths being piecewise constant, with a finite number of jumps, this a rare case in functional data analysis in which the trajectories are not supposed to be continuous and can be observed exhaustively. The approach is illustrated on a data set of sensory perceptions, considering different gustometer-controlled stimuli experiments. We also show how it can be easily extended to analyze experiments, such as temporal check-all-that-apply, in which two states or more can be observed at the same time.
Keywords : categorical functional data, continuous time categorical processes, dimension reduction, discontinuous trajectories, sensometrics, temporal dominance of sensations, temporal check all that apply
@@ 1 Introduction
A lot of attention has been paid in the statistical literature over the last decades to develop tools dedicated to the analysis and modeling of functional data, considering random trajectories defined on an interval $[0,T]$ and taking values in $\mathbb{R}$, at each $t\in[0,T]$ (see for example Gertheiss et al., (2024) and Koner and Staicu, (2023) for recent reviews of the literature and Ramsay and Silverman, (2005) for a seminal reference on functional data analysis). Much less attention has been given to the case in which the trajectories take values in a finite set of elements that are not necessarily numbers, that is to say when $\mathbb{R}$ is replaced by the finite set $\mathcal{S}=\{S_{1},\ldots,S_{q}\}$, with cardinality $q$.
However, there are many examples in which the statistical units of interest are samples of a continuous time random categorical process : in a pioneer work on demographic studies, Deville, (1982) extended the notion of multiple correspondence analysis to continuous time correspondence analysis in order to analyze the time evolution of the ”marital status” of a sample of women over the period $[1962,1975]$. The trajectories related to the marital status take values in a state space $\mathcal{S}$ with $q=4$ states (divorced, married, single, widowed). Recent examples of statistical analysis of individual categorical trajectories are found in food science, a domain in which it is of great interest to get information about the temporal perception of aliments to understand the perception mechanisms. Of particular interest is the Temporal Dominance of Sensations (TDS) approach which consists in choosing sequentially attributes, among a list composed of $q$ predefined items, describing a food product over tasting (see Pineau et al., (2009)). The chosen states correspond to the most striking perception at a given time and the results of TDS experiments, after time normalization, can be represented via barplots as in Figure 1. The technique developed by Deville, (1982), also called qualitative harmonic analysis (see Deville and Saporta, (1980)) or categorical functional data analysis (CFDA) is now available in R, package cfda (see Preda et al., (2021)) and can be used to analyze such kind of data. Even if the CFDA approach can be very powerful by encoding categorical trajectories into a sequence of real components (see Peltier et al., (2023) for an illustration in sensory analysis), the interpretation of the results in terms of individual trajectories is often delicate.
@@
Statistical approaches based on Markov processes and their extensions (see Lindsey, (2012) for an overview and Limnios and Oprişan, (2001) for an introduction to semi-Markov processes) can be useful to fit the law of the trajectories at the population level and to provide a simple representation of the dynamics via the graph of transitions between states. Considering parametric distributions for the sojourn time in the different states also allows to deal with maximum likelihood estimation techniques, permitting to build two-sample tests to compare two populations (e.g two food products or two categories of consumers (see Cardot and Frascolla, (2024)) as well as model-based clustering, considering mixtures of semi-Markov processes (see Cardot et al., (2019)). A major drawback of this approach is that it is not well suited for analyzing data at the individual level, and the Markovian assumption is often too simplistic to properly fit real data. Additionally, it is difficult to apply when more than one state can be present simultaneously (see the TCATA experiment described in Section 4.2), as this requires drastically increasing the number of system states to $2^{q}$ to account for all possible combinations.
@@
One could consider such trajectories as compositional data evolving over time (see Aitchison, (1983) for a seminal paper and Greenacre, (2021) for a recent review on compositional data analysis). A major difficulty to deal with compositional data approaches in our framework is due to the fact that we deal with individual trajectories so that, at each instant $t$, q1q-1 individual observations, among $X_{1}(t),\ldots,X_{q}(t)$, have value 0. In other words, at each instant $t$, all the observed units are on the vertices of the simplex, so that usual logarithmic transforms cannot apply directly.
@@
Our approach is based on the optimal approximation, according to the $L^{2}$ distance, to such 0-1 trajectories in a small dimension functional space. This leads us to consider the multivariate functional principal component analysis (MFPCA, see Chiou et al., (2014) or Happ and Greven, (2018)) of the multivariate functional vector $\mathbf{X}=(X_{1},\ldots,X_{q})$. We show that it leads both to principal components that can be simply interpreted in terms of variations around the mean probability curve related to each state, and powerful tools able to reduce effectively the dimension of categorical functional data in a finite dimension vector space. All codes in the R language are available on Github, https://github.com/Chemosens/ExternalCode/tree/main/MFPCAWithCategoricalTrajectories.
2 Notations and mathematical framework
@@
To describe continuous time categorical trajectories, as those drawn in Figure 1, we consider the random element $Y=\{Y(t),\ t\in[0,T]\}$, with $Y(t)\in\mathcal{S}=\{S_{1},\ldots,S_{q}\}$ for all $t\in[0,T]$. We denote by $p_{j}(t)=\mathbb{P}\left[Y(t)=S_{j}\right]$, the probability of being in state $S_{j}$ at time $t$ and, for $(s,t)\in[0,T]\times[0,T]$ and $(j,\ell)\in\{1,\ldots,q\}\times\{1,\ldots,q\}$, the joint probability
$p_{j\ell}(s,t)=\mathbb{P}\left[Y(s)=S_{j}\mbox{ and }Y(t)=S_{\ell}\right].$
We associate to $Y$, $q$ trajectories $X_{j}$, $j=1,\ldots,q$, related to the occurrence of state $S_{j}$, and defined as follows, $X_{j}(t)=1$ if $Y(t)=S_{j}$ and 0 otherwise. In other words, $X_{j}(t)=\mathbf{1}_{\{Y(t)=S_{j}\}}$, for all $t\in[0,T]$, where $\mathbf{1}_{\{.\}}$ is the indicator function. We clearly have
$\mathbb{E}[X_{j}(t)]$ $=\mathbb{E}[\mathbf{1}_{\{Y(t)=S_{j}\}}]$
$=p_{j}(t)$
and
$\mathbb{E}[X_{j}(s)X_{\ell}(t)]$ $=\mathbb{E}[\mathbf{1}_{\{Y(s)=S_{j}\}}\mathbf{1}_{\{Y(t)=S_{\ell}}\}]$
@@
ensuring that the trajectories $X_{j},\ j=1,\ldots,q$ are continuous in probability. Hypothesis $\mathbf{H}_{1}$ prevents them to have too many jumps or to have jumps that occur at the same time point $t_{0}$ with non null probability. Define the covariance functions $\gamma_{j}(s,t)=\mathbb{C}ov(X_{j}(s),X_{j}(t))$ and, for $\ell\neq j$, $\gamma_{j\ell}(s,t)=\mathbb{C}ov(X_{j}(s),X_{\ell}(t))$. We remark that $\gamma_{j}(t,t)=p_{j}(t)(1-p_{j}(t))$ and $\gamma_{j\ell}(s,t)=p_{j\ell}(s,t)-p_{j}(s)p_{\ell}(t)$, so that the terms out of the diagonal, when $s\neq t$, can be related to a departure from independence.
@@
of Proposition 2.1
Remarking that $\sup_{t\in[0,T]}|X_{j}(t)|\leq 1$ almost surely, we deduce, by Theorem 1.3.6 in Serfling, (1980), that the trajectories are also continuous in the $L^{2}$ sense (or mean square continuous) when $\mathbf{H}_{1}$ is true, that is to say
@@
The continuity $p_{j}$ and $\gamma_{jj}$ is a consequence of Theorem 7.3.2 in Hsing and Eubank, (2015) which states that the mean $p_{j}$ and covariance functions $\gamma_{jj}$ are continuous if and only if $X_{j}$ is mean-square continuous. To prove the continuity of $\gamma_{j\ell}$, we note that for $(s,t)$ and $(s^{\prime},t^{\prime})$ in $0,T]\times[0,T]$, we get thanks to the Cauchy-Schwarz inequality,
@@
Note that the binary trajectories $X_{j}$ take values in the Hilbert space $L^{2}[0,T]$ equipped with the usual inner product, denoted by $\langle.,.\rangle$, and norm $\|.\|$. Denote by $\Gamma_{j\ell}:L^{2}[0,T]\to L^{2}[0,T]$ the cross-covariance operator between $X_{j}$ and $X_{\ell}$. It is the integral operator with kernel function $\gamma_{j\ell}(s,t)$,
$\Gamma_{j\ell}f(s)=\int_{0}^{T}\left(p_{j\ell}(s,t)-p_{j}(s)p_{\ell}(t)\right)f(t)dt.$ (2)
Considering now simultaneously $X_{1},\ldots,X_{q}$, we denote by $\mathbf{X}=(X_{1},\ldots,X_{q})$ the random vector of functions, that takes values in $\mathbb{H}=L^{2}[0,T]\times\cdots\times L^{2}[0,T]$, the Hilbert space of $q$ dimensional vectors of functions in $L^{2}[0,T]$, equipped with the inner product
$\langle\mathbf{f},\mathbf{g}\rangle_{\mathbb{H}}$ $=\sum_{j=1}^{q}w_{j}\langle f_{j},g_{j}\rangle.$ (3)
The weights $w_{j}$ that can be chosen by the statistician (see Remark 2 below and Section 4) are strictly positive. We denote by $\mathbf{p}=(p_{1},\ldots,p_{q})\in\mathbb{H}$ the expectation of $\mathbf{X}$, $\mathbf{p}=\mathbb{E}[\boldsymbol{X}]$, and by $\boldsymbol{\Gamma}:\mathbb{H}\to\mathbb{H}$ the covariance operator. It satisfies, $\forall\boldsymbol{\phi}=(\phi_{1},\ldots,\phi_{q})\in\mathbb{H}$,
$\boldsymbol{\Gamma}\boldsymbol{\phi}(s)$ $=\mathbb{E}\left[\langle\mathbf{X}-\mathbf{p},\boldsymbol{\phi}\rangle_{\mathbb{H}}\left(\mathbf{X}(s)-\mathbf{p}(s)\right)\right],$ (4)
$=\sum_{\ell=1}^{q}w_{\ell}\begin{pmatrix}\Gamma_{1\ell}\phi_{\ell}(s)\\ \vdots\\ \Gamma_{q\ell}\phi_{\ell}(s)\end{pmatrix}$ (5)
We denote by $\boldsymbol{\gamma}(s,t)$ the corresponding (weighted) multivariate covariance function, with elements $\left(w_{\ell}\gamma_{j\ell}(s,t)\right)_{j,\ell}$. We deduce from Proposition 2.1 and the multivariate version of Mercer’s theorem (see Chiou et al., (2014) or Happ and Greven, (2018)) that there exists a set of orthonormal basis functions $\boldsymbol{\phi}_{r}(s)=(\phi_{r1}(s),\ldots,\phi_{rq}(s))$ in $\mathbb{H}$, $r=1,2,\ldots$ and corresponding eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq 0$ such that
$\boldsymbol{\gamma}(s,t)=\sum_{r\geq 1}\lambda_{r}\boldsymbol{\phi}_{r}(s)\boldsymbol{\phi}_{r}(t)^{\top}.$
This leads to the following expansion of the covariance functions $\gamma_{j\ell}(s,t)$,
$p_{j\ell}(s,t)-p_{j}(s)p_{\ell}(t)$ $=\sum_{r\geq 1}\lambda_{r}\phi_{rj}(s)\phi_{r\ell}(t).$ (6)
It is not difficult to show that operator $\boldsymbol{\Gamma}$, defined in (4), is the integral operator with kernel function $\boldsymbol{\gamma}(s,t)$ and $\boldsymbol{\Gamma}\boldsymbol{\phi}_{r}=\lambda_{r}\boldsymbol{\phi}_{r}$. Furthermore, $\sum_{r\geq 1}\lambda_{r}=\mathbb{E}\left[\|\mathbf{X}-\mathbf{p}\|_{\mathbb{H}}^{2}\right]\leq T\max_{j}w_{j}$.
$\widetilde{\mathbf{X}}_{k}(t)$ $=\mathbf{p}+\sum_{r=1}^{k}\langle\mathbf{X}-\mathbf{p},\boldsymbol{\phi}_{r}\rangle_{\mathbb{H}}\boldsymbol{\phi}_{r}(t).$ (7)
@@
The continuous time extension of correspondence analysis, named qualitative harmonic analysis and developed by Deville, (1982) and Deville and Saporta, (1980) is based on the eigen decomposition of another integral operator, with a purpose that is not to expand the trajectories themselves in an ”optimal” way but to relate the states $S_{j}$ to numerical values at each instant $t$. More precisely, the aim is to find an optimal encoding function $\varphi:\mathcal{S}\times[0,1]\to\mathbb{R}$ minimizing the following criterion
$\int_{0}^{T}\!\!\!\!\int_{0}^{T}\mathbb{E}\left[\left(\varphi(Y(t),t)-\varphi(Y(s),s)\right)^{2}\right]dsdt$
subject to identifiability constraints $\mathbb{E}\left[\varphi(Y_{t},t)\right]=0$ for all $t\in[0,1]$ and unit variance,
$\int_{0}^{T}\mathbb{E}\left[\varphi(Y(t),t)^{2}\right]dt=1$. A solution $\varphi(x,t)$ satisfies, for all $x\in\mathcal{S}$, the integral operator equation (see equation (42) in Deville, (1982)),
@@
Departure from independence is evaluated via the ratio $\frac{p_{j\ell}(t,s)}{p_{j}(t)p_{\ell}(s)}$ which is equal to one in case of independence. Denoting by $\widetilde{\lambda}_{1}\geq\widetilde{\lambda}_{2}\geq\ldots\geq 0$ the sequence eigenvalues of operator equation (8) and by $\boldsymbol{\varphi}_{i}(t)=(\varphi_{i}(S_{1},t),\ldots,\varphi_{i}(S_{q},t))$ the eigenfunction related to $\widetilde{\lambda}_{i}$, we get a Mercer type expansion of the joint probabilities (see equation (43) in Deville, (1982)),
@@
If the statistician wants to give the same importance to all the states, one reasonable option consists in choosing weights
@@
$=\left(\int_{0}^{T}p_{j}(t)\left(1-p_{j}(t)\right)dt\right)^{-1}$ (10)

### v4

# Statistical description and dimension reduction of continuous time categorical trajectories with multivariate functional principal components

@@ Abstract
Getting tools that allow simple representations and comparisons of a set of categorical trajectories is of major interest for statisticians. Without loosing any information, we associate to each state a binary random indicator function, taking values in $\{0,1\}$, and turn the problem of statistical description of the categorical trajectories into a multivariate functional principal components analysis. This viewpoint encompasses experimental frameworks where two or more states can be observed simultaneously. The sample paths being piecewise constant, with a finite number of jumps, this a rare case in functional data analysis in which the trajectories are not supposed to be continuous and can be observed exhaustively. Under the weak hypothesis assuming only continuity in probability of the 010-1 trajectories, the means and the (multivariate) covariance function are continuous and have interpretations in terms of departure from independence of the joint probabilities. Considering a functional data point of view, we show that the binary trajectories, which are right-continuous functions with left-hand limits, can be seen as random elements in the Hilbert space of square integrable functions. The multivariate functional principal components are simple to interpret and we show that we can get consistent estimators of the mean trajectories and the covariance functions under weak regularity assumptions. The ability of the approach to represent categorical trajectories in a small dimension space is illustrated on a data set of sensory perceptions, considering different gustometer-controlled stimuli experiments.
Keywords : categorical functional data, continuous time categorical processes, dimension reduction, discontinuous trajectories, jump processes, sensometrics, telegraph process, temporal dominance of sensations, temporal check all that apply.
@@ 1 Introduction
A lot of attention has been paid in the statistical literature over the last decades to develop tools dedicated to the analysis and modeling of functional data, considering random trajectories defined on an interval $[0,T]$ and taking values in $\mathbb{R}$, at each $t\in[0,T]$ (see for example Gertheiss et al.,, 2024 and Koner and Staicu,, 2023 for recent reviews of the literature and Ramsay and Silverman,, 2005 for a seminal reference on functional data analysis). Much less attention has been given to the case in which the trajectories take values in a finite set of elements that are not necessarily numbers, that is to say when $\mathbb{R}$ is replaced by the finite set $\mathcal{S}=\{S_{1},\ldots,S_{q}\}$, with cardinality $q$.
However, there are many examples in which the statistical units of interest are samples of a continuous time random categorical process : in a pioneer work on demographic studies, Deville, (1982) extended the notion of multiple correspondence analysis to continuous time correspondence analysis in order to analyze the time evolution of the “marital status” of a sample of women over the period $[1962,1975]$. The trajectories related to the marital status take values in a state space $\mathcal{S}$ with $q=4$ states (divorced, married, single, widowed). Recent examples of statistical analysis of individual categorical trajectories are found in food science, a domain in which it is of great interest to get information about the temporal perception of aliments to understand the perception mechanisms. Of particular interest is the Temporal Dominance of Sensations (TDS) approach which consists in choosing sequentially attributes, among a list composed of $q$ predefined items, describing a food product over tasting (see Pineau et al.,, 2009). The chosen states correspond to the most striking perception at a given time and the results of TDS experiments, after time normalization, can be represented via barplots as in Figure 1. The technique developed by Deville,, 1982, also called qualitative harmonic analysis (see Deville and Saporta,, 1980) or categorical functional data analysis (CFDA) is now available in R, package cfda (see Preda et al.,, 2021) and can be used to analyze such kind of data. Even if the CFDA approach can be very powerful by encoding categorical trajectories into a sequence of real components (see Peltier et al.,, 2023 for an illustration in sensory analysis), the interpretation of the results in terms of individual trajectories is often delicate since it is not the purpose of the method.
@@
Statistical approaches based on Markov processes and their extensions (see Lindsey,, 2012 for an overview and Limnios and Oprişan,, 2001 for an introduction to semi-Markov processes) can be useful to fit the law of the trajectories at the population level and to provide a simple representation of the dynamics via the graph of transitions between states. Considering parametric distributions for the sojourn time in the different states also allows to deal with maximum likelihood estimation techniques, permitting to build two-sample tests to compare two populations, e.g two food products or two categories of consumers (see Cardot and Frascolla,, 2024) as well as model-based clustering, considering mixtures of semi-Markov processes (see Cardot et al.,, 2019). A major drawback of this approach is that it is not well suited for analyzing data at the individual level, and the Markovian assumption is often too simplistic to properly fit real data. Additionally, it is difficult to apply when more than one state can be present simultaneously (see for example the TCATA experiment described in Castura et al.,, 2016 and in the Supplementary Material), as this requires drastically increasing the number of system states to $2^{q}$ to account for all possible combinations.
@@
One could consider such trajectories as compositional data evolving over time (see Aitchison,, 1983 for a seminal paper and Greenacre,, 2021 for a recent review on compositional data analysis). A major difficulty to deal with compositional data approaches in our framework is due to the fact that we deal with individual trajectories so that, at each instant $t$, q1q-1 individual observations, among $X_{1}(t),\ldots,X_{q}(t)$, have value 0. In other words, at each instant $t$, all the observed units are on the vertices of the simplex, so that usual logarithmic transforms cannot apply directly.
The approach presented in this work is an extension of Cardot and Peltier, (2025), giving a more detailed mathematical description of the properties of the trajectories and estimators of the mean and the covariance. This work encompasses both the TDS and TCATA experimental frameworks, the latter allowing for the simultaneous observation of two or more states.
Under the assumption that the trajectories are continuous in probability and have almost surely a finite number of jumps we get that such random elements are jointly measurable (see Hsing and Eubank,, 2015, Chapter 7) and we deduce that they can be seen as random elements in a multidimensional Hilbert space of square integrable functions. This leads us to consider the multivariate functional principal component analysis (MFPCA) of the multidimensional functional vector $\mathbf{X}=(X_{1},\ldots,X_{q})$. When having a sample of trajectories at hand, consistency results are obtained both from a functional data and continuous time stochastic process point of view.
@@
We show on the example presented in Figure 1 that it leads both to principal components that can be simply interpreted in terms of variations around the mean probability curve related to each state, and powerful tools able to reduce effectively the dimension of categorical functional data in a finite dimension vector space. We also present a comparison, on this case study, with the CFDA approach.
All the codes in the R language (see R Core Team,, 2024) are available on Github, https://github.com/Chemosens/ExternalCode/tree/main/MFPCAWithCategoricalTrajectories. Additional numerical experiments, including weighted MFPCA, TCATA analysis, and jump time measurements corrupted by noise, along with an example based on the telegraph process, where explicit calculations are derived, are provided in the Supplementary Material.
2 Mathematical framework
Consider a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and suppose that the collection of random variables $Y(t)$, taking values in $\mathcal{S}=\{S_{1},\ldots,S_{q}\}$, are well defined for each $t$ in $[0,T]$. We can consider, in a equivalent way, the collection of $q$-dimensional random vectors $\mathbf{X}(t)=(X_{1}(t),\ldots,X_{q}(t))$, where $X_{j}(t)$, $j=1,\ldots,q$, are binary stochastic processes, related to the occurrence of state $S_{j}$ at time $t$, and defined as follows. For all $\omega\in\Omega$, $X_{j}(t,\omega)=1$ if $Y(t,\omega)=S_{j}$ and 0 otherwise. In other words, $X_{j}(t)=\mathbf{1}_{\{Y(t)=S_{j}\}}$, for all $t\in[0,T]$, where $\mathbf{1}_{\{.\}}$ is the indicator function.
As in Billingsley, (1999) or Hsing and Eubank, (2015), it is possible consider two distinct point of views to describe mathematically the continuous time categorical trajectories drawn in Figure 1. On the one hand we can consider the continuous time stochastic process point of view and study the properties of $\mathbf{X}(t)$ defined for each $t$ in $[0,T]$. On the other hand we can consider a description based on a functional point of view by studying the random element $\mathbf{X}(.,\omega)=\{(X_{1}(t,\omega),\ldots,X_{q}(t,\omega)),\ t\in[0,T]\}$, taking values in a function space, for each $\omega\in\Omega$.
Note that in TDS experiments, where exactly one state is selected at time $t$, the multivariate stochastic process $\mathbf{X}(t)$ follows a multinomial distribution with $n=1$ trial and a success probability vector $\mathbf{p}(t)=(p_{1}(t),\ldots,p_{q}(t))$. In TCATA experiments, however, the values of $Y(t,\omega)$ are no longer single elements of $\mathcal{S}$ but subsets of $\mathcal{S}$. We define $X_{j}(t)=\mathbf{1}_{\{Y(t)\ni S_{j}\}}$, which takes the value 1 if the state $S_{j}$ is observed at time $t$, that is to say $S_{j}\in Y(t,\omega)$, and 0 otherwise. We have $\sum_{j=1}^{q}X_{j}(t)\in\{0,\ldots,q\}$ and the joint distribution of $\mathbf{X}(t)$ is no longer multinomial. Nevertheless, the marginal distribution of $X_{j}(t)$ remains Bernoulli with success probability denoted by $p_{j}(t)$.
Unless explicitly stated otherwise, we will assume from now on that $\mathbf{X}(t)$ is the result of a TDS or TCATA type experiment.
2.1 The continuous time stochastic process point of view
@@
We denote by $p_{j}(t)=\mathbb{P}\left[Y(t)\ni S_{j}\right]$, the probability of observing state $S_{j}$ at time $t$ and, for $(s,t)\in[0,T]\times[0,T]$ and $(j,\ell)\in\{1,\ldots,q\}\times\{1,\ldots,q\}$, the joint probability
$p_{j\ell}(s,t)=\mathbb{P}\left[Y(s)\ni S_{j}\mbox{ and }Y(t)\ni S_{\ell}\right].$
We clearly have
$\mathbb{E}[X_{j}(t)]$ $=\mathbb{E}[\mathbf{1}_{\{Y(t)\ni S_{j}\}}]$
$=p_{j}(t).$
Consider two distinct time points $(s,t)$, we have
$\mathbb{E}[X_{j}(s)X_{\ell}(t)]$ $=\mathbb{E}[\mathbf{1}_{\{Y(s)\ni S_{j}\}}\mathbf{1}_{\{Y(t)\ni S_{\ell}\}}]$
@@
ensuring that the trajectories of $X_{j},\ j=1,\ldots,q$ are continuous in probability. Hypothesis $\mathbf{H}_{1}$ prevents them to have too many jumps or to have jumps that occur at a same time point $t_{0}\in[0,T]$ with non null probability. Define the covariance functions $\gamma_{j}(s,t)=\mathbb{C}ov(X_{j}(s),X_{j}(t))$ and, for $\ell\neq j$, $\gamma_{j\ell}(s,t)=\mathbb{C}ov(X_{j}(s),X_{\ell}(t))$. We remark that $\gamma_{j}(t,t)=p_{j}(t)(1-p_{j}(t))$ and $\gamma_{j\ell}(s,t)=p_{j\ell}(s,t)-p_{j}(s)p_{\ell}(t)$, so that the terms out of the diagonal, when $s\neq t$, can be related to a departure from independence.
@@
Remarking that $\sup_{t\in[0,T]}|X_{j}(t)|\leq q$ almost surely, we deduce, by Theorem 1.3.6 in Serfling, (1980), that the trajectories are also continuous in the $L^{2}$ sense (or mean square continuous) when $\mathbf{H}_{1}$ is true, that is to say
@@
The continuity of $p_{j}$ and $\gamma_{jj}$ is then a consequence of Theorem 7.3.2 in Hsing and Eubank, (2015) which states that the mean $p_{j}$ and covariance functions $\gamma_{jj}$ are continuous if and only if $X_{j}$ is mean-square continuous. To prove the continuity of $\gamma_{j\ell}$, we note that for $(s,t)$ and $(s^{\prime},t^{\prime})$ in $0,T]\times[0,T]$, we get thanks to the Cauchy-Schwarz inequality,
@@
Although stochastic processes with values in $\{0,1\}$ can exhibit very smooth mean functions (see Appendix C in the Supplementary Material), their trajectories, being discontinuous due to jumps of size 1, are not mean square differentiable. This lack of regularity translates into non-differentiability of the covariance function near the diagonal $(t,t)$, as discussed in Chapter 11, Loève, (1978).
As in Hsing and Eubank, (2015), Chapter 7, we define the cross-covariance operator between $X_{j}$ and $X_{\ell}$ as the integral operator $\Gamma_{j\ell}:L^{2}[0,T]\to L^{2}[0,T]$, such that for all $f\in L^{2}[0,T]$,
$\Gamma_{j\ell}f(s)=\int_{0}^{T}\left(p_{j\ell}(s,t)-p_{j}(s)p_{\ell}(t)\right)f(t)dt,\quad\forall s\in[0,T].$ (2)
Note that since $p_{j\ell}(s,t)-p_{j}(s)p_{\ell}(t)$ is a continuous function, $\Gamma_{j\ell}f$ is also a continuous function (see Lemma 4.6.1 in Hsing and Eubank,, 2015).
Let us introduce some notation. Consider the separable Hilbert space $L^{2}[0,T]$ equipped with the usual inner product, denoted by $\langle\cdot,\cdot\rangle$, and norm $\|\cdot\|$. We further define the vector space $\mathbb{H}=L^{2}([0,T],\mathbb{R}^{q})$ as the separable Hilbert space of functions $\mathbf{f}=(f_{1},\ldots,f_{q})$ defined on $[0,T]$, with values in $\mathbb{R}^{q}$ such that $\|\mathbf{f}\|_{\mathbb{H}}^{2}=\sum_{j=1}^{q}\|f_{j}\|^{2}<\infty$. For any $\mathbf{u}=(u_{1},\dots,u_{q})$ and $\mathbf{v}=(v_{1},\dots,v_{q})$ in $\mathbb{H}$, the inner product is defined component-wise as follows,
$\langle\mathbf{u},\mathbf{v}\rangle_{\mathbb{H}}=\sum_{j=1}^{q}\langle u_{j},v_{j}\rangle.$ (3)
We denote by $\boldsymbol{\Gamma}:\mathbb{H}\to\mathbb{H}$ the multivariate integral covariance operator as follows. For all $\mathbf{f}=(f_{1},\ldots,f_{q})\in\mathbb{H}$,
$\boldsymbol{\Gamma}\mathbf{f}$ $=\sum_{\ell=1}^{q}\begin{pmatrix}\Gamma_{1\ell}f_{\ell}\\ \vdots\\ \Gamma_{q\ell}f_{\ell}\end{pmatrix}$. (4)
The operator $\boldsymbol{\Gamma}$ can be seen as the integral operator with multivariate kernel function $\boldsymbol{\gamma}(s,t)$, with elements $\left(\gamma_{j\ell}(s,t)\right)_{j,\ell}$. We deduce from Proposition 2.1 and the multivariate version of Mercer’s theorem (see Chiou et al.,, 2014 or Happ and Greven,, 2018) that there exists a set of continuous and orthonormal basis functions $\boldsymbol{\phi}_{r}=(\phi_{r1},\ldots,\phi_{rq})\in\mathbb{H}$, $r=1,2,\ldots$ and corresponding eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq 0$ such that
$\boldsymbol{\gamma}(s,t)$ $=\sum_{r\geq 1}\lambda_{r}\boldsymbol{\phi}_{r}(s)\boldsymbol{\phi}_{r}(t)^{\top},\quad\forall(s,t)\in[0,T]\times[0,T],$ (5)
where the sum converges absolutely and uniformly on $0,T]\times[0,T]$. This leads to the following expansion of the covariance functions $\gamma_{j\ell}(s,t)$,
$p_{j\ell}(s,t)-p_{j}(s)p_{\ell}(t)$ $=\sum_{r\geq 1}\lambda_{r}\phi_{rj}(s)\phi_{r\ell}(t),\quad\forall(s,t)\in[0,T]\times[0,T],$ (6)
where the infinite sum converges uniformly on $0,T]\times[0,T]$.
@@
The continuous time extension of correspondence analysis, named qualitative harmonic analysis and developed by Deville, (1982) and Deville and Saporta, (1980) is based on the eigendecomposition of another integral operator, with a purpose that is not to expand the trajectories themselves in an “optimal” way but to relate the states $S_{j}$ to numerical values at each instant $t$. More precisely, the aim is to find an optimal encoding function $\varphi:\mathcal{S}\times[0,T]\to\mathbb{R}$ minimizing the following criterion
$\int_{0}^{T}\!\!\!\!\int_{0}^{T}\mathbb{E}\left[\left(\varphi(Y(t),t)-\varphi(Y(s),s)\right)^{2}\right]dsdt,$ (7)
subject to identifiability constraints $\mathbb{E}\left[\varphi(Y_{t},t)\right]=0$ for all $t\in[0,T]$ and unit variance,
$\int_{0}^{T}\mathbb{E}\left[\varphi(Y(t),t)^{2}\right]dt=1$. A solution $\varphi(x,t)$ satisfies, for all $x\in\mathcal{S}$, the integral operator equation (see equation (42) in Deville,, 1982),
@@
Departure from independence is evaluated via the ratio $\frac{p_{j\ell}(t,s)}{p_{j}(t)p_{\ell}(s)}$ which is equal to one in case of independence. Denoting by $\widetilde{\lambda}_{1}\geq\widetilde{\lambda}_{2}\geq\ldots\geq 0$ the sequence eigenvalues of operator equation (8) and by $\boldsymbol{\varphi}_{i}(t)=(\varphi_{i}(S_{1},t),\ldots,\varphi_{i}(S_{q},t))$ the eigenfunction related to $\widetilde{\lambda}_{i}$, we get a Mercer type expansion of the joint probabilities (see equation (43) in Deville,, 1982),
@@
Note that it is possible to introduce positive weights $w_{1},\ldots,w_{q}$ in the inner product defined in (3):
$\langle\mathbf{u},\mathbf{v}\rangle_{\mathbb{H}}=\sum_{j=1}^{q}w_{j}\langle u_{j},v_{j}\rangle.$
These weights can be chosen by the statistician in order to give the same importance to all the states, by considering for example
@@
$=\left(\int_{0}^{T}p_{j}(t)\left(1-p_{j}(t)\right)dt\right)^{-1},$ (10)